One of the engineers has been reviewing the data and thinks humidity may be a factor that is affecting the packaging process. The engineer tracked the humidity levels for a random sample of cereal packages over 30 days along with the humidity level on that day during production.  

1. Calculate the correlation coefficient, r.

2. Interpret the correlation coefficient, r. Be sure to address both the direction and strength of the association between package weights and humidity level.

3. What next steps would you recommend the company to take based on these results? Be specific, creative, and logical.

Weight Humidity
19.987 19
19.997 19
20.02 20
20.117 22
20.21 21
20.222 22
20.226 22
20.258 23
20.321 23
20.334 24
20.396 24
20.412 24
20.434 25
20.468 25
20.527 26
20.541 25
20.541 26
20.554 27
20.555 29
20.556 28
20.611 31
20.616 30
20.622 31
20.634 32
20.64 31
20.691 32
20.721 33
20.773 34
20.816 36
21.002 38
Answer #1

1. X Values
? = 613.802
Mean = 20.46
?(X - Mx)2 = SSx = 1.827

Y Values
? = 802
Mean = 26.733
?(Y - My)2 = SSy = 757.867

X and Y Combined
N = 30
?(X - Mx)(Y - My) = 35.622

R Calculation
r = ?((X - My)(Y - Mx)) / ?((SSx)(SSy))

r = 35.622 / ?((1.827)(757.867)) = 0.9573

2. Y2 0.95732 0.9164

Which means 91.64% of variation in humidity is explained by thickness

3. Based on this observation we should suggest that in order to have an effective packaging process kindly take care that humidity is been maintained accordingly